OpenStudy (anonymous):

Find the difference quotient of f, that is find (f(x+h) - f(x))/ (h) for the function f(x)= x^2 + 5x - 1

OpenStudy (turingtest):

at what point are you stuck?

OpenStudy (anonymous):

I just really don't understand what the question is asking and what I am supposed to do. Do you want me to rewrite the equation to make it look better?

OpenStudy (turingtest):

no I understand it fine thanks :) what is f(x+h) ? to find it, just replace every x in f(x) with x+h what do you get?

OpenStudy (anonymous):

\[f(x) = x ^{2} + 2xh +h ^{2} + 5h + 5x -1\]

OpenStudy (turingtest):

yes :) so what is f(x+h)-f(x)?

OpenStudy (anonymous):

@Brybro where did u get the 5h?

OpenStudy (anonymous):

So do I just take that equation I just posted and subtract from the original f(x) equation right?

OpenStudy (anonymous):

would it be.... x^2+2xh+h^2+4x-1?

OpenStudy (turingtest):

right

OpenStudy (turingtest):

@dsrnation no

OpenStudy (anonymous):

@dsrnation give me a sec. I will post my work soon

OpenStudy (anonymous):

@TuringTest can you check my answer after im done?

OpenStudy (turingtest):

sure thing

OpenStudy (anonymous):

\[f(x) = h ^{2} + 2xh + 10x + 5h - 2\]

OpenStudy (turingtest):

careful what you write, that is not f(x) but rather f(x+h)-f(x)

OpenStudy (turingtest):

actually it's wrong anyway :P

OpenStudy (anonymous):

x^2+2xh+h^2=4x=3h-1?

OpenStudy (anonymous):

meant for the = tp be +

OpenStudy (turingtest):

\[f(x+h)=x^2+2xh+h^2+5x+5h-1\]\[f(x)=x^2+5x-1\]\[f(x+h)-f(x)=x^2+2xh+h^2+5x+5h-1-(x^2+5x-1)\]

OpenStudy (turingtest):

\[f(x+h)-f(x)=2xh+h^2+5h\]

OpenStudy (anonymous):

Oh i forgot to distribute the - sign. That makes since

OpenStudy (turingtest):

yep, so now the only thing missing is the "over h" so divide this all by h and that is the difference quotient

OpenStudy (anonymous):

i know this isnt my problem but thanks!!! i just started learning these!

OpenStudy (turingtest):

no prob :)

OpenStudy (anonymous):

So the answer is 2x + h + 5?

OpenStudy (turingtest):

yes, nice job

OpenStudy (anonymous):

Thanks so much. I have a test tomorrow and I'm not the best at math. Think you could hel me with one more question for now?

OpenStudy (turingtest):

I'm a bit busy, but post it separately and I'll help as much as I can

OpenStudy (anonymous):

No problem. Thanks